OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3 These are the options : a. y = -3x - 13 b. y = -3x + 11 c. y = -3x + 13 d. y = -3x + 1

OpenStudy (anonymous):

Please Help?

jigglypuff314 (jigglypuff314):

slope-intercept form: y = mx + b when m = slope and b = y-intercept given slope was -3 so plug in as m = -3 to get y = -3x + b follow so far? :)

OpenStudy (anonymous):

Not at all lol i'm confused

OpenStudy (anonymous):

Is it y= -3x + 11 ?

OpenStudy (anonymous):

First part

jigglypuff314 (jigglypuff314):

there is a format called "the slope intercept form" that looks like y = mx + b where m and b are numbers and m would represent the slope

OpenStudy (anonymous):

So y = -3x + 11 ?

jigglypuff314 (jigglypuff314):

not quite :) y = -3x + b to find the value of that ^ number temporarily plug in point (2, -5) as (x, y) into the equation so solve for " b " for -5 = -3(2) + b

OpenStudy (anonymous):

My bad i think is y = -3x - 11 ?

jigglypuff314 (jigglypuff314):

-5 = -3(2) + b -5 = -6 + b then add 6 to both sides so b = -5 + 6 = ?

OpenStudy (anonymous):

is it y = 3x -11.. Honetsly im completely lost :/

OpenStudy (anonymous):

i'm not good at math

jigglypuff314 (jigglypuff314):

it's alright :) just do -5 + 6 = ?

OpenStudy (anonymous):

-5 + 6 = 1

jigglypuff314 (jigglypuff314):

yup :) so we got from -5 = -3(2) + b to --> b = 1 so y = -3x + b and since b = 1 y = -3x + 1

OpenStudy (anonymous):

Got it :] Thanks .